OpenStudy (anonymous):

What is the sign of the product? (–0.7)(0.7)(–2.9)(–3.6)(–7.1)

OpenStudy (anonymous):

A. positive B. negative

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Positive

OpenStudy (anonymous):

thx @ObeyFCS

OpenStudy (anonymous):

neg*pos=neg*neg=pos*neg=neg*neg=pos